# Identifying Sparse Treatment Effects

### v1

@@ 6.1 Simulations
We compare the performance of the baseline and the proposed methods with pre-treatment covariates adjustment, as presented in Section 5, through a simulation study. In scenarios where the data generation process involves independent outcomes and a constant treatment effect without pre-treatment covariates, both the baseline and proposed approaches demonstrate similar performance. Additional details can be found in the Appendix C. However, in a more realistic and complicated setting as follows, our proposed method outperforms the baseline method in most cases.
We generate data as follows. First, we randomly sample the model parameters as
@@
for $j=1,\dots,p$, where $p$ is the outcome dimension. Next, for the generated model parameters $\beta$ and $\delta$, we randomly generate $n$ samples in the following manner:
for $i=1,\dots,n$ and $j=1,\dots,p$. Note that there are non-zero treatment effects on $S_{\tau}=\{Y_{1},\dots,Y_{s_{\tau}}\}$, where $s_{\tau}$ is a pre-specified sparsity index in the simulation process.
We evaluate the performance of the proposed method relative to the baseline method, both with adjustment for pre-treatment covariates, based on how well the ranking recovers $S_{\tau}$. The ranking is obtained as follows: For the baseline method, individual outcome dimensions are ranked based on their treatment effect sizes. For the proposed method, we consider regularization paths for a fine grid of $\lambda$ values and track a corresponding sequence of active sets (the set with non-zero coefficients) of increasing size. The ranking is determined by the order in which outcome dimensions enter the Lasso path. The recovery rate is defined as follows: For a given $s$, denote $S(s)$ as the selected subset of outcome dimensions of size $s$ obtained from applying each method with pre-treatment covariates adjustment. Then, for each method, the recovery rate of $S_{\tau}$ is $|S(s)\cap S_{\tau}|/s$ for $1\leq s\leq s_{\tau}$.
The recovery rates comparing the proposed to baseline approaches (with pre-treatment covariates adjustment) for different sample sizes ($n$), probabilities of being treated ($\pi$), and magnitudes of treatment effects ($\alpha$) are given in Figure 3. We set the outcome dimension $p=100$, the number of pre-treatment covariates $m=30$, and the sparsity index $s_{\tau}=5$. Note that in simulations, we use the elastic net penalty [Zou and Hastie, 2005] instead of the Lasso penalty as the elastic net is known to encourage a grouping effect: correlated predictors tend to be in or out of the model together. We see that our proposed method outperforms baseline methods in most scenarios, especially when sample sizes are small and signals are weak. When sample sizes and signals are large, the proposed method tends to identify a subset of $S_{\tau}$ more quickly than baseline methods, but performs slightly worse in discovering the full $S_{\tau}$.
Figure 3: Recovery rates averaged over 1000 replicates for the baseline and proposed approaches with pre-treatment covariates adjustment for $n=100,200,300,400,500$ and $\alpha=0.2,0.4,0.6,0.8$ for $\pi=0.3$ (Above) and $\pi=0.5$ (Below).
@@ 6.2 Semi-Synthetic Data
In this section, we demonstrate the performance of our method using semi-synthetic data motivated from the time series data example introduced in Section 1. Following the protocol discussed in Section 1, we curate a dataset with $n=5,000$ samples of the average 24-hour period for a week of real CGM data before and after a potential treatment. For each sample $i$, let the transformed glucose traces be $\text{glucose}_{i}\in\mathbb{R}^{288\times 2=576}$, where 288 represents 24 hours of data recorded at 5 minute intervals averaged over the 7 days in a given week and we concatenate this 24-hour period average representation for Week 1 and Week 2. An example of such a transformed CGM trace is illustrated in Figure 1.
Generalizing the procedure of Section 1, we generate semi-synthetic data by simulating a treatment effect that occurs within a random two-hour time interval of the day. For example, the treatment effect might occur following lunch, between 12pm and 2pm. Given a randomly selected two-hour interval $I$ in which the treatment effect may manifest, a synthetic constant treatment effect of magnitude $\alpha$ is applied as
@@
if sample $i$ is treated and $t\in I$. As before, we randomize treatment assignment with probability of 0.5 of a sample receiving a treatment.
@@
In contrast to using pre-defined windows, our proposed approach allows for exploration of more refined windows while achieving superior power. In the first stage, we randomly partition the data set into two splits. Within the first split, we aim to locate the treatment effect and consider multiple resolutions for the representation. We denote 6 four-hour windows as level 0, 12 two-hour windows as level 1, and 24 one-hour windows as level 2. We perform subset selection with our proposed method where we set the pre-treament covariates $X$ to be TIR of considered windows of the pre-treatment week (Week 1) and $Y$ to be TIR of considered windows of the post-treatment week (Week 2). We perform subset selection for each level and compute the RSS for the selected subset within the level. The level and corresponding selected subset with the smallest RSS is chosen. Moving to the second split, we estimate the treatment effect utilizing the selected subset of the chosen level. We apply the multi-sample splitting procedure of Meinshausen et al. [2008] to our entire process to reduce the dependence on the random data split; see Algorithm 2.
The entire process, involving the generation of a random two-hour treatment effect and the application of methods introduced above, is repeated 1000 times. The powers of the considered methods from these 1000 iterations are presented in Figure 4. We see that our method achieves better power than the default method of using pre-specified representations and multiple testing.
Figure 4: Power plots for the semi-synthetic glucose data comparing our proposed method (blue) to a default method of looking at a pre-specified representation of 4-hour windows (red) or 2-hour windows (yellow) and then applying multiple testing.
@@ 7 Discussion
This paper addresses the problem of causal inference from experimental data with high-dimensional outcomes. Our emphasis is on the development of a practical, easily deployed method that can identify the sparse subset of outcome representations most likely to contain the treatment effect given a large pool of outcome representations. We introduce a sparse regression approach for subset selection, where we employ a $\ell_{1}$-regularized weighted linear regression with carefully chosen weights. We provide theoretical results showing that our method can recover the subset of outcome representations consistently. In simulated and semi-synthetic experiments, we empirically demonstrate that our method outperforms the baseline methods, especially when the signal-to-noise ratio is relatively weak.
Looking ahead, a key challenge lies in the sample-splitting framework. Multiple correction testing may experience decreased power due to a potentially large correction factor. However, methods based on sample splitting may suffer from reduced power as they utilize only half of the data for treatment effect estimation. Compared to the baseline (individual ranking based) subset selection method, our method performs better in low sample-size settings, suggesting the potential for allocating more samples to the second data split for the actual treatment effect estimation. Nonetheless, the overall process necessitates sufficient sample sizes. There are many real-world examples where the primary challenge is obtaining adequate sample sizes. For those cases, alternative approaches beyond sample splitting may be required. However, there are also many scenarios where sample sizes are sufficient, but the size of the treatment effect is small; our procedures is particularly well-suited to such scenarios. More generally, in the setting where scientists believe the treatment effect is sparse and are interested in selecting the outcome representations in which the treatment effect potentially manifests, our approach provides practical guidance on how to do so.
8 Acknowledgement
The authors would like to thank Kevin Guo, James Yang, Samir Khan, Yu Wang, and Tim Morrison for their insightful comments and helpful feedback. The authors would also like to thank Johannes Ferstad for assistance with the data set used in the experiments above.
@@ B.5 Proof of Lemma 1
Appendix C Additional Simulations
C.1 Without Pre-treament Covarites
@@
C.2 With Partially-Observable Pre-treatment Covariates
When pre-treament covariates are partially observable under the simulation model in Section 6.1, the sparsity assumption of the inverse covariance matrix in Section 5.2 is violated. In the following, we add simulation results where we only observe a subset of pre-treament covariates; we only observe $20$ out of $m=30$. The results are presented in Figure 6. We see that the proposed approach still mostly outperforms the baseline approach, especially when sample sizes are small and signals are weak.
Figure 6: Recovery rates averaged over 1000 replicates for the baseline and proposed approaches with pre-treatment covariates adjustment when pre-treament covariates are partially observable for $n=100,200,300,400,500$ and $\alpha=0.2,0.4,0.6,0.8$ for $\pi=0.3$ (Above) and $\pi=0.5$ (Below).

### v2

@@ 6.1 Simulations
We compare the ranking performance and power of the baseline and the proposed methods with pre-treatment covariate adjustment, as presented in Section 5, through a simulation study. In scenarios where the data generation process involves independent outcomes and a constant treatment effect without pre-treatment covariates, both the baseline and proposed approaches demonstrate similar performance. Additional details can be found in the Appendix D. Therefore, we focus on scenarios where pre-treatment covariates explain some of the variance in outcomes.
We generate $n$ samples in the following manner:
for $i=1,\dots,n$ and $j=1,\dots,p$, where $p$ is the outcome dimension. The model parameters are first randomly sampled as
@@
for $j=1,\dots,p$. Note that there are non-zero treatment effects on $S_{\tau}=\{Y_{1},\dots,Y_{s_{\tau}}\}$, where $s_{\tau}$ is a pre-specified sparsity index in the simulation process. We set the outcome dimension $p=500$, the number of pre-treatment covariates $m=50$, and the sparsity index $s_{\tau}=5$.
We evaluate the performance of the proposed method relative to the baseline method, both with adjustment for pre-treatment covariates, based on how well the ranking recovers $S_{\tau}$. The ranking is obtained as follows: For the baseline method, individual outcome dimensions are ranked based on their treatment effect sizes. For the proposed method, we consider regularization paths for a fine grid of $\lambda$ values and track a corresponding sequence of active sets (the set with non-zero coefficients) of increasing size. The recovery rate is defined as follows: For a given $s$, denote $S(s)$ as the selected subset of outcome dimensions of size $s$ obtained from applying each method with pre-treatment covariates adjustment. Then, for each method, the recovery rate of $S_{\tau}$ is defined as $|S(s)\cap S_{\tau}|/s$ for $1\leq s\leq s_{\tau}$.
Additionally, we evaluate statistical power of both methods. A second data set with a sample size of 500 was generated from the same data-generating process. For each selected subset of size $s$ from the first data set with a sample size of $n$, we evaluate power on the second data set for each method.
The recovery rates and powers comparing the proposed to baseline approaches (with pre-treatment covariates adjustment) for different sample sizes ($n$), probabilities of being treated ($\pi$), and magnitudes of treatment effects ($\alpha$) are given in Figure 3 and 4. Note that in our simulations, we use both the Lasso penalty and the elastic net penalty [Zou and Hastie, 2005]. The elastic net is known for encouraging a grouping effect, where correlated predictors tend to be included or excluded together. We see that our proposed method outperforms baseline methods in most scenarios, especially when sample sizes are small and signals are weak, with the elastic net achieving even better performance than the Lasso. When sample sizes and signals are large, the proposed method tends to identify a subset of $S_{\tau}$ more quickly than baseline methods, but performs slightly worse in discovering the full $S_{\tau}$.
Figure 3: Recovery rates and powers averaged over 1000 replicates for the baseline and proposed approaches with pre-treatment covariates adjustment for $n=200,300,400,500$ and $\alpha=0.4,0.6,0.8,1$ for $\pi=0.3$.
Figure 4: Recovery rates and powers averaged over 1000 replicates for the baseline and proposed approaches with pre-treatment covariates adjustment for $n=200,300,400,500$ and $\alpha=0.4,0.6,0.8,1$ for $\pi=0.7$.
@@ 6.2 Semi-Synthetic Data
In this section, we demonstrate the performance of our method using semi-synthetic data motivated from the time series data example introduced in Section 1. Following the protocol discussed in Section 1, we curate a dataset with $n=5,000$ samples of the average 24-hour period for a week of real CGM data before and after a potential treatment. Specifically, we randomly sampled 10,000 two-week CGM traces from the CGM time-series data of 149 patients collected from the ongoing 4T study at Stanford’s Lucille Packard Children’s Hospital. While these traces are not fully independent—since multiple traces can come from the same patient—we assume independence for our semi-synthetic analysis. From the 10,000 traces, we retained the 5,000 with the fewest missing values.
For each sample $i$, let the transformed glucose traces be $\text{glucose}_{i}\in\mathbb{R}^{288\times 2=576}$, where 288 represents 24 hours of data recorded at 5 minute intervals averaged over the 7 days in a given week and we concatenate this 24-hour period average representation for Week 1 and Week 2. An example of such a transformed CGM trace is illustrated in Figure 1. Generalizing the procedure of Section 1, we generate semi-synthetic data by simulating a treatment effect that occurs within a random two-hour time interval of the day. The two-hour time interval is randomly selected across the 1000 repeats of the procedure. For example, the treatment effect might occur following lunch, between 12pm and 2pm. Given a randomly selected two-hour interval $I$ in which the treatment effect may manifest, a synthetic constant treatment effect of magnitude $\alpha$ is applied as
@@
if sample $i$ is treated and $t\in I$. We randomize treatment assignment with probability of 0.5 of a sample receiving a treatment after matching samples with the nbpmatching package in R.
@@
In contrast to using a fixed length of windows, our proposed approach allows for exploration of more refined windows while achieving better power. In the first stage, we randomly partition the data set into two splits. Within the first split, we aim to locate the treatment effect and consider multiple resolutions for windows. We denote 6 four-hour windows as level 0, 12 two-hour windows as level 1, and 24 one-hour windows as level 2. We perform subset selection with our proposed method where we set the pre-treament covariates $X$ to be TIR of considered windows of the pre-treatment week (Week 1) and $Y$ to be TIR of considered windows of the post-treatment week (Week 2). We perform subset selection for each level and compute the RSS as described in Section 5.1 for the selected subset within the level. The level and corresponding selected subset with the smallest RSS is chosen. Moving to the second split, we estimate the treatment effect utilizing the selected subset of the chosen level. We apply the multi-sample splitting procedure of Meinshausen et al. [2008] to our entire process to reduce the dependence on the random data split; see Algorithm 2.
The entire process, involving the generation of a random two-hour treatment effect and the application of methods introduced above, is repeated 1000 times. The powers of the considered methods from these 1000 iterations are presented in Figure 5. We see that our method achieves better power than the default method of using fixed-sized windows and multiple testing. Our method performs better for several reasons: (1) the treatment effect is sparse, and our approach efficiently captures this sparsity by allowing for flexibility in the identification of the treatment effect windows; (2) when the fixed-window size is too large compared to the actual treatment effect, the signal is diluted, leading to reduced power in the baseline method; and (3) even when the fixed-window size is as narrow as the actual sparse treatment effects, the baseline method still suffers from power loss due to the multiple testing corrections applied to each window. Our method, however, can identify the treatment effect window from the first stage and proceed without substantial multiple testing corrections. If multiple windows are selected in the first stage in our method, a multiple testing correction factor is still applied, but it is generally smaller than that required by the baseline method.
Figure 5: Power plots for the semi-synthetic glucose data comparing our proposed method (blue) to a default method of looking at pre-specified 4-hour windows (red) or 2-hour windows (yellow) and then applying multiple testing.
@@ 7 Discussion
This paper addresses the problem of causal inference from experimental data with high-dimensional outcomes. Our emphasis is on the development of a practical, easily deployed method that can identify the sparse subset of outcome summaries most likely to contain the treatment effect. We first cast the problem as one of maximizing an objective based on the Hotelling’s $T^{2}$ test statistic; to get to computational tractability, we then equate this objective to performing subset selection within a weighted linear regression setting with carefully chosen weights. For this task, we deploy $\ell_{1}$ regularization. We provide theoretical results showing that our method can recover the appropriate subset of outcome summaries consistently. In simulated and semi-synthetic experiments, we empirically demonstrate that our method outperforms the baseline methods, especially when the signal-to-noise ratio is relatively weak.
A key challenge lies when one wants to employ our method with sample-splitting. Sample-splitting based methods may suffer from reduced power as they utilize only half of the data for the treatment effect estimation. Compared to the baseline (individual-ranking-based) subset selection method, our method performs better in low sample-size settings, suggesting the potential for allocating more samples to the second data split for the actual treatment effect estimation. Nonetheless, the overall process necessitates sufficient sample sizes. In settings where the primary challenge is obtaining adequate sample sizes, alternative approaches beyond sample splitting may be required. That said, our procedure is particularly adept if the dimensionality of the outcome summaries far exceeds sample sizes even when sample sizes are sufficient for the second-stage inference (assuming sparsity in the first stage); this is a common scenario in analysis of gene expression data. In clinical settings, which are our main motivation, we often begin with a small pilot or POC study, followed by a much larger clinical trial. In these cases, where the second stage involves a much larger sample size, our method can be effectively applied to select relevant outcome summaries and estimate treatment effects.
8 Acknowledgments
The authors would like to thank Dominik Rothenhausler, Kevin Guo, James Yang, Samir Khan, Yu Wang, and Tim Morrison for their helpful feedback. The authors would also like to thank Johannes Ferstad for assistance with the dataset used in the experiments above. This work was supported in part by the National Science Foundation Grant 2205084; the Helmsley Charitable Trust; NIH/NIDDK Grant R18DK122422; AFOSR Grant FA9550-21-1-0397, ONR Grant N00014-22-1-2110; the Stanford Institute for Human-Centered AI (HAI); and the Stanford Maternal and Child Health Research Institute. EBF is a Chan Zuckerberg Biohub - San Francisco Investigator. YJ is supported by the Stanford Data Science Scholars Program.
@@ B.5 Proof of Lemma 1
Appendix C Multi Sample-Splitting
Extension to Multi-Sample Splitting.
Methods based on sample splitting are known to be sensitive to the arbitrary choice of a one-time random split of the data [Meinshausen et al., 2008]. Therefore, it is recommended to employ multiple splits and aggregate p-values from each split to improve robustness of results. In Algorithm 2, we present an algorithm where we apply the multi-sample splitting results from Meinshausen et al. [2008] to our proposed method. Note that one consequence of using multiple splits is that while we obtain an aggregated p-value for testing the null hypothesis, we no longer obtain an estimated subset localizing the treatment effect (since each split produces one such subset). Aggregating the subsets remains an interesting direction for future work.
Input: $(T_{i},X_{i},Y_{i})_{i=1}^{n}$, $\gamma$
Output: valid p-values under $H_{0}:\tau=0$
1 for b = 1, …, B do
3       Calculate p-value:
  1. (a)
    (P-values with multiple testing) The p-value for each $Y_{j}$ is
    $p_{j}^{(b)}$ $=1\quad\text{for }j\notin\hat{S}_{\tau}^{(b)},$
    where $\Phi$ is the cdf of the standard Gaussian distribution.
  2. (b)
    (P-value with Hotelling $T^{2}$ statistic) The p-value is
Aggregate p-values as
Algorithm 2 Multi-Split Algorithm
Appendix D Additional Simulations
D.1 Without Pre-treament Covarites
@@
D.2 With Partially-Observable Pre-treatment Covariates
When pre-treament covariates are partially observable under the simulation model in Section 6.1, the sparsity assumption of the inverse covariance matrix in Section 5.2 is violated. In the following, we add simulation results where we only observe a subset of pre-treament covariates; we only observe $40$ out of $m=50$. The results are presented in Figure 7 and 8. We see that the proposed approach still mostly outperforms the baseline approach, especially when sample sizes are small and signals are weak.
Figure 7: Recovery rates and powers averaged over 1000 replicates for the baseline and proposed approaches with pre-treatment covariates adjustment when pre-treament covariates are partially observable for $n=200,300,400,500$ and $\alpha=0.4,0.6,0.8,1$ for $\pi=0.3$.
Figure 8: Recovery rates and powers averaged over 1000 replicates for the baseline and proposed approaches with pre-treatment covariates adjustment when pre-treament covariates are partially observable for $n=200,300,400,500$ and $\alpha=0.4,0.6,0.8,1$ for $\pi=0.7$.